OpenStudy (anonymous):

What is the 17th term of the arithmetic sequence 18, 22, 26, … ?

OpenStudy (anonymous):

an = a + (n-1)d

OpenStudy (anonymous):

d = t2 - t1 = 22 - 18 = 4

OpenStudy (anonymous):

a = 18 n = 17

OpenStudy (anonymous):

First calculate common difference.... as @Yahoo! did

OpenStudy (anonymous):

u see i have given

OpenStudy (anonymous):

@Katiebae .... did u understand

OpenStudy (anonymous):

Kind of

OpenStudy (anonymous):

wat??

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

where didnt u understand

OpenStudy (anonymous):

So, a1=18 a2=22 a3=26

OpenStudy (anonymous):

did u understand this...... @Katiebae

OpenStudy (anonymous):

@Katiebae .............. why do u always relpy very late

OpenStudy (anonymous):

I was making breakfast for my brother

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

Ok........ did u understand?

OpenStudy (anonymous):

@Katiebae I am tired of waiting

OpenStudy (anonymous):

Yes I do

OpenStudy (jiteshmeghwal9):

hey @Katiebae look carefully to the question one number is adding with 4 to gt next number

OpenStudy (anonymous):

OK (:

OpenStudy (jiteshmeghwal9):

like 18 18+4=22 22+4=26 take this term to the 17th sequence

razor99 (razor99):

hey i could help

razor99 (razor99):

@Katiebae

razor99 (razor99):

u there

razor99 (razor99):